• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name ransome

    Log entry time 15:53:34 on November 6,1999

    Entry number 29938

    keyword=Shift summary

    Started with gamma,d generally smooth running except for
    ROC 14 problems.
    11:12 went to e,d
    beam was unstable, low current

    12:40 beam off for re-tune of injector, went to hydrogen target
    12:55 beam back at 30 uA
    Apparently there was some mis-steering at the injector,
    problem now fixed.

    Target didn't react well, took a few minutes to stablize
    13:30 go back to gamma d
    14:05 stop data taking to prepare for Moller run
    Moller to end of shift